Solve
Divide 84 into two parts that are in the ratio of 3:4.

Answer by Cromlix(4381)

Hi there,
Divide 84 into two parts that are in the ratio of 3:4.
3 + 4 = 7
3/7 x 84 = 36
4/7 x 84 = 48
Total......84
Hope this helps :-)